OpenStudy (anonymous):

A software company is releasing one of their products on a CD. The manufacturer charges a $5,500 setup fee and $1.00 per CD. Approximately how many CDs must be produced so that the total cost per CD is $3.50?

OpenStudy (anonymous):

Total cost C = x + 5500 where x = number of CDs Find the point at which C/x = 3.5

OpenStudy (anonymous):

i still dont understand

geerky42 (geerky42):

A software company had to purchase a setup fee and CDs, which cost C, which is equal to x + 5500, where x is a number of CDs. Company wants to sell CDs at $3.50 each to make up for what it purchased. How many CDs does company have to sell? Just like Rav said, total cost C = 1.00x + 5,500.00 where x = number of CDs. You have to find the point at which C/1.00x = 3.5 C/x = 3.5 Multiply both sides by x. C = 3.5x C = x + 5500 3.5x = x + 5500 \(\vdots\) Is this clear? Can you find x?